### v1

# Expanding the Boundaries of Vision Prior Knowledge in Multi-modal Large Language Models

Abstract
Does the prior knowledge of the vision encoder constrain the capability boundary of Multi-modal Large Language Models (MLLMs)? While most existing research treats MLLMs as unified systems optimized through end-to-end training, the impact of vision encoder’s prior knowledge is seldom investigated. In this work, we introduce a novel metric, $Rank_{e}$, to quantify the effect of the vision encoder’s prior knowledge on MLLM performance. Our analysis reveals a positive correlation between prior knowledge and MLLM performance. Moreover, we find that domain-specific fine-tuning using solely end-to-end visual question answering (VQA) data is insufficient—particularly for entities with low inherent visual prior knowledge. To address this issue, we propose VisPRE (Vision Prior Remediation), a two-stage training framework that explicitly incorporates prior knowledge at the vision encoder level. Experimental results demonstrate that augmenting vision encoder’s prior knowledge substantially boosts the visual understanding capabilities of MLLMs, offering a novel and effective strategy for improving performance, especially in scenarios involving uncommon visual entities.
11footnotetext: Equal contribution.
1 Introduction
Multi-modal Large Language Models have emerged as a rapidly growing area of research. Combining the powerful capabilities of Large Language Models with the ability to process visual input, MLLMs excel in tasks such as image understanding, VQA [1], image captioning, and visual instruction following. The development of models like GPT-4o[30] , GPT-4V[29], and Claude-3.5[3] have demonstrated remarkable proficiency in advanced multi-modal understanding. Besides, open-source models like LLaVA[26, 25, 18] series, Qwen2-VL[38], and InternVL2[9, 8] are making significant strides, bridging gap in the field.
Figure 1: Knowledge quadrants of a MLLM. “Vision known” indicates that the vision encoder recognises the entity in the image, while “Language known” indicates that the language model possesses relevant information about the entity. Only when both vision and language are “known” can the MLLM achieve accurate comprehension and response generation.
Figure 2: Left: Current MLLM performance vs. vision prior knowledge. Current MLLMs demonstrate positive correlation between vision prior knowledge and overall performance. Right: “Vision Known” and “Vision Not Known” Entities. (1) For “vision known entities”, the vision encoder contains sufficient prior knowledge, enabling MLLM answers correctly; (2) For “vision not known entities”, insufficient visual knowledge leads to MLLM failure. We propose the $Rank_{e}$ metric to quantify vision encoder’s prior knowledge about specific entities, along with a two-stage training framework to enhance encoder knowledge, expanding MLLM’s performance boundaries.
A pivotal challenge in advancing MLLMs is forging a seamless and robust alignment between vision and language. One effective approach involves integrating an off-the-shelf external vision encoder with a language model using a modality conversion module [2, 21, 23, 42, 13, 4, 26, 19], which we refer to as the modular approach. Compared to monolithic multi-modal approach [34, 27, 5, 41], which built from scratch using multi-modal data, modular approach is more data-efficient and achieves comparable performance. Despite these advantages, the modular approach still faces challenges, as the vision and language components are trained separately from distinct data distributions, leading to an inherent misalignment in their knowledge. To illustrate the importance of knowledge alignment, we present a knowledge quadrant diagram in Fig. 1 , with the horizontal axis representing the knowledge held by language model and the vertical axis representing the knowledge held by the vision encoder. Only when both components possess necessary knowledge (in the “Vision known & Language known” quadrant) can the multi-modal model accurately handle complex cross-modal tasks [22, 10]. Misalignment in knowledge from either vision or language side introduces limitations to the model’s capabilities, making it essential to bridge this gap to enhance the performance of multi-modal models. Many existing studies focus on addressing knowledge misalignment from the language perspective, expanding from “Vision known & Language not known” to “Vision known & Language known”. Some studies[6, 17] enhance language model knowledge with external documents related to images, while CVLM[24] trains a “Visual Knowledge Aligner” module to enrich text-based knowledge associated with images. However, as a crucial component of MLLM [12], the vision encoder also possesses varying prior knowledge about the real world, such as entities, textures and causality [32, 7]. But the impact of this vision prior knowledge on MLLM capabilities remains unexplored, leading to a natural question: How does vision prior knowledge affect MLLM’s capability? In this paper, we attempt to answer this question by investigating the following research questions:
  • Q1: How to measure prior knowledge in vision encoders?
  • Q2: Does vision prior knowledge constrain MLLM?
  • Q3: How to transcend vision prior knowledge limits?
To address these questions, we introduce $Rank_{e}$ to quantify the vision encoder’s prior knowledge. Through experiments with various model combinations, we reveal a positive correlation between MLLM performance and visual prior knowledge. Fig. 2 (left) demonstrates the relationship between current MLLM performance and vision prior. Furthermore, we find that direct finetuning with end-to-end VQA data is insufficient for improving low prior entities. Fig. 2 (right) illustrates the knowledge misalignment on low prior entities. To overcome this limitation, we propose a two-stage training framework that injects vision prior knowledge into the vision encoder, resulting in significant improvements in MLLM performance. Through in-depth analysis, we demonstrate that the limitations of direct finetuning stem from the differences between next-word prediction loss and contrastive loss. In summary, our main contributions are:
  • We introduce the $Rank_{e}$ metric to quantify a vision encoder’s prior knowledge, revealing a positive correlation between MLLM performance and the encoder’s embedded visual knowledge.
  • Our analysis shows that domain-specific finetuning with only end-to-end VQA data proves insufficient, particularly for entities with low vision prior knowledge.
  • We propose a two-stage training framework VisPRE (Vision Prior Remediation) that injects prior knowledge at the vision encoder level, significantly enhancing MLLM performance, especially for entities with vision prior knowledge.
2 Vision Prior Measurement
Vision encoders are typically trained on extremely large-scale data (from 400 million to 10 billion samples[36]), often with undisclosed data (e.g. OpenAI CLIP[33]), making direct evaluation of vision priors from training data infeasible. Therefore, to answer Q1, we shift our focus to evaluating observable behavioral evidence - specifically, how effectively these encoders recognize visual entities. We thus propose the $Rank_{e}$ metric, which quantifies an encoder’s vision prior knowledge for a given entity $e$.
In this section, we begin by describing the modality alignment process in modular MLLMs, then formulating a definition of vision prior knowledge. Finally, we introduce the $Rank_{e}$ metric to quantify this knowledge.
Modern modular MLLMs establish cross-modal understanding through an alignment process that maps visual information to textual representations. Formally, given an input text prompt $T_{A}$ and target image $I_{B}$, the alignment function $\mathcal{F}$ of an MLLM generates textual description that approximate the image content:
[…]
Here, $\hat{T}_{B}$ represents the generated text that preserves the semantic content of $I_{B}$. Building upon the Platonic representation hypothesis [16], we posit that cross-modal alignment occurs through a shared latent space $\mathcal{Z}$. This allows us to decompose the $P(T|I_{B})$ as:
$P(T|I_{B})=\sum_{z\in\mathcal{Z}}\underbrace{P_{\text{vision}}(z|I_{B})}_{\text{Vision prior}}\cdot P_{\text{align}}(T|z,I_{B})$ (2)
The latent representation $z$ serves as an intermediary that connects the visual and textual domains. While $P_{\text{align}}(T|z,I_{B})$ reflects the MLLM’s ability to convert latent representation $z$ into textual output $T$, $P_{\text{vision}}(z|I_{B})$ represents the vision encoder’s capability to transform image $I_{B}$ into an appropriate latent representation. $P_{\text{vision}}(z|I_{B})$ constitutes what we define as vision prior knowledge—the encoder’s pre-existing understanding of visual entities encoded in its parameters.
To quantify the inherent vision prior $P_{\text{vision}}(z|I_{B})$, we discretize the continuous latent space $\mathcal{Z}$ into a set of entity-specific latent representations. For a given image $I_{B}$, we approximate $P(z|I_{B})$ by evaluating the probability that the vision encoder correctly identifies entity within $I_{B}$. To achieve this, we propose the $Rank_{e}$ metric, which measures how well the encoder identifies a target entity $e$ from visual inputs, thereby evaluating the vision encoder’s inherent prior knowledge. As shown in Fig. 3, for an entity $e$, $Rank_{e}$ is computed as follows:
  1. 1.
    Similarity scoring: For an image $I_{e}$ containing entity $e$, compute the image-text similarity score $s_{j}=\phi(I_{e},T_{j})$ using the vision encoder and its corresponding text encoder, where $\{T_{1},...,T_{n}\}$ are textual descriptions of $n$ candidate entities.
  2. 2.
    Ranking: Rank the entities in descending order based on their similarity scores $\{s_{j}\}_{j=1}^{n}$, and record the position of the target entity $e$ as $Rank_{e}$. If multiple images $\{I_{e}^{(1)},...,I_{e}^{(m)}\}$ are available for single entity $e$, compute $Rank_{e}$ for each image separately and take the average:
    $\text{Rank}_{e}=\frac{1}{m}\sum_{i=1}^{m}rank\big{(}\phi(I_{e}^{(i)},T_{e})\big{)}.$ (3)
    where $rank(\phi(I_{e},T_{e}))$ denotes the position of $\phi(I_{e},T_{e}))$ in ordered $\{s_{j}\}_{j=1}^{n}$. Lower $Rank_{e}$ values indicate stronger visual prior knowledge, with optimal performance when $\text{Rank}_{e}=1$.
Figure 3: Illustration of metric Ranke. For a target entity $e$, we compute cross-modal similarity scores between its vision representations (extracted by vision encoder) and text representations of all candidate entities (extracted by corresponding text encoder). The rank of entity $e$ among these candidates defines its $Rank_{e}$. In this example, while Image A depicts Entity A, entity A achieves 4th-highest similarity score, resulting in $Rank_{e}$ = 4.
3 Experiments
In this section, we explore the three proposed research questions. In Section 3.1, we describe the overall experimental setup. In Section 3.2, we verify the relationship between MLLM and the prior knowledge of its vision encoder. From Section 3.3 to Section 3.4, we show the insufficient of end-to-end finetuning and propose a training framework to transcend vision prior knowledge limits.
3.1 Experiment Setting
Models. To systematically examine the impact of the vision encoder’s prior knowledge on MLLM performance across different vision encoders and base LLM combinations, we train nine MLLMs from scratch based on an encoder-projector-LLM architecture. For the vision encoder, we utilize widely adopted encoders in mainstream MLLMs, including OpenAI ViT-L-14[33], SigLIP ViT-SO-14[40], and DFN ViT-H-14[15]. For the base LLM, we select the LLaVA-1.5 language model, Vicuna-7B-v1.5[11], and the recent open-source models, Llama-3.1-Instruct-7B[14] and Qwen-2.5-Instruct-7B[35]. This results in a total of nine encoder-language model combinations.
Datasets. To evaluate MLLMs under different vision priors, we require a VQA dataset that meets two conditions: (1) it provides entity annotations covering a wide range of prior knowledge—from extremely rare to very common entities; (2) it includes entity-centric visual questions and answers for MLLM performance assessment. Here, rare entities refer to those that appear infrequently or not at all in the vision encoder’s training data, making them difficult for the vision encoder to recognize accurately. The Encyclopedia-VQA[28] dataset fulfills both requirements. With extensive entity annotations covering up to 16.7k entity categories, it captures both common and rare entities and poses a hard challenge for MLLMs with its knowledge-based VQA questions.
Training. We trained the model on 8×A800 GPUs. We follow the LLaVA[26] training process and made a few modifications. We first pre-train model on LLaVA pretrain data to obtain a projector aligned with selected vision encoder. To reduce computational cost, we then sampled 10% of the LLaVA instruction tuning dataset and used it in the finetuning strategy for our subsequent experiments.
Metrics and Evaluation. We use Llama-3.1-70B[14] to judge model responses, denoted as a function $g(\cdot)$ that takes the question, entity, ground truth answer, and model output as input, returning true if the answer is correct. Using this, we define entity accuracy $\text{Acc}_{e}$ for each entity $e$ as the fraction of correct responses among all related questions:
[…]
where $N_{e}$ is the number of questions for entity $e$, $y_{i}$ is the ground truth answer and other question information, and $\hat{y}_{i}$ is the model’s output. The overall dataset accuracy $\text{Acc}_{\text{macro}}$ is calculated as the macro-average of all entity accuracies. Details of the evaluation configurations are in Appendix B.
Figure 4: MLLM Performance distribution across different Ranke intervals. Performance of all MLLMs decreases as $Rank_{e}$ increases across three encoder configurations. The Vicuna-CLIP model shows an 87% performance drop from $50000<Rank_{e}<$ 5000 to $Ranke>3000$, indicating correlation between performance and vision prior knowledge. This relationship is non-linear with a critical threshold. We marked this threshold by a vertical line in the figure—green on the left indicating sufficient prior knowledge for reasoning, and red on the right showing insufficient knowledge causing sharp performance decline.
3.2 Vision Prior Constrains MLLM Performance
To investigate Q2: the impact of visual priors on MLLMs, we first categorize entities into two types: those “vision encoder knows” and those “vision encoder doesn’t know” then observe MLLM performance across both categories. Through our proposed $Rank_{e}$ metric, we measure the vision encoder’s knowledge of entities in Encyclopedia-VQA, where a lower $Rank_{e}$ indicates greater knowledge. For MLLM performance, we test accuracy in answering entity-related questions in Encyclopedia-VQA.
Our study aims at addressing knowledge misalignment where MLLM capabilities are limited by the vision encoder. Therefore, we retain only cases where the LLM component possesses adequate entity knowledge, regardless of the vision encoder’s knowledge. Specifically, we prompt the MLLM with “This is {entity name}” rather than the actual image; if the MLLM answers correctly, we retain this case. Additionally, we discovered a number of cases where MLLMs provide correct answer without image description or actual image. We attribute this to the MLLM’s dependency on question format [17]. We eliminated this subset from our analysis. Fig. 4 illustrates the relationship between MLLM accuracy and $Rank_{e}$
Finding 1: MLLM performance correlates positively with vision prior knowledge. As shown in Fig. 4, across all three encoder choices, MLLM performance consistently declines as entity $Rank_{e}$ increases. For the CLIP encoder, from the interval $5000<Rank_{e}<$ 500 to $Rank_{e}>3000$, Vicuna’s performance drops by 87%, Llama3.1’s by 100%, and Qwen-2.5’s by 21%. In SigLIP encoder experiments, overall performance declines by about 50% across all three models from the leftmost to the rightmost interval, while for the DFN encoder, the decline reaches 100%.
Notably, CLIP-Vicuna MLLM does not exhibit a significant performance decline until $Rank_{e}$ reaches 3000. The phenomenon is also observed in the SigLIP and DFN configurations. This threshold effect suggests that the positive correlation between vision prior knowledge and MLLM performance is not strictly linear, but rather exhibits a mutation beyond a critical point. We posit that this stems from the vision encoder holding a known status for entities below a certain $Rank_{e}$ threshold, meaning it can still provide sufficient prior knowledge for the MLLM to answer entity-related questions. Once $Rank_{e}$ exceeds this threshold, the vision encoder no longer provides adequate prior knowledge, resulting in a sharp drop in MLLM performance. Considering that LLM part of MLLM possesses adequate knowledge about all entities here, so it is the vision encoder of MLLM that constrains the overall performance on entities beyond the threshold.
3.3 Limitations of End-to-end Finetuning
To investigate Q3: How to transcend vision prior knowledge limits, we implement a typical solution as our baseline—finetuning MLLMs on end-to-end domain-specific VQA data. Following established MLLM finetuning approaches [26, 25], we freeze the vision encoder parameters and only tune the LLM component. This setup enables the LLM parameters to compensate for limitations in vision prior knowledge.
Vision Encoder LLM Number of (Q, A) pairs Number of entities
   Train    Test
OpenAI ViT-L-14 Vicuna-7B    1877    531 90
Llama3.1-8B    2305    624 106
Qwen2.5-7B    2345    645 109
SigLIP ViT-SO-14 Vicuna-7B    2290    615 106
Llama3.1-8B    2669    717 123
Qwen2.5-7B    2614    705 118
DFN ViT-H-14 Vicuna-7B    1914    531 90
Llama3.1-8B    2339    615 105
Qwen2.5-7B    2291    618 105
Table 1: Dataset Statistics. We report the number of (question, answer) pairs entities for dataset split for different encoder-language model combinations. Note that the training set and testset of each dataset contain the same entities.
We construct our finetuning dataset from Encyclopedia-VQA. Following method in Section 3.2, we retain questions that MLLMs answer correctly when prompted with “This is {entity name}” instead of the actual image. After calculating $Rank_{e}$ across the dataset, we observe significantly different $Rank_{e}$ distributions for different encoders. To mitigate this distributional variance, we sample entities with more uniform rank distributions for validation. Each subset is then divided into training and test sets, ensuring they contain the same entities with no question duplication. Dataset statistics are presented in Tab. 1, with detailed construction methodology in Appendix A.
Successful knowledge-based VQA requires three MLLM capabilities: (1) recognizing entities in images; (2) possessing relevant knowledge about these entities; and (3) utilizing this knowledge to answer questions. Since we have already ensured the LLM component possesses adequate entity knowledge, MLLM performance can be enhanced through two different approaches: (1) improving entity recognition and (2) optimizing knowledge utilization.
To explore these approaches, we develop two distinct types of finetuning data: (1) Perception-tuning data, where we transform original Encyclopedia-VQA questions into perception-focused queries such as What is this image about? and (2) Knowledge-tuning data, which preserves the original questions from Encyclopedia-VQA. Detailed construction methodologies for both datasets are provided in Appendix A.
Figure 5: Perception-tuning and Knowledge-tuning performance improvement compared to Zero-shot. Perception-tuning provides limited improvement to model performance across different $Rank_{e}$ values, with a significant drop observed in Qwen-2.5’s performance within the range of $Rank_{e}>3000$. For Knowledge-tuning, we observe that in the range $Rank_{e}>3000$, both Qwen-2.5 and Llama-3.1 models finetuned with Knowledge-tuning data experience a notable decline in performance.
Finding 2: Domain-specific finetuning with only end-to-end VQA data is insufficient, particularly for entities with low visual prior knowledge. Fig. 5 illustrates the accuracy improvements of Perception-tuning and Knowledge-tuning models compared to Zero-shot baselines under CLIP encoder configuration. As shown in figure (a), after Perception-tuning, Qwen-2.5 performance decreased in the $Rank_{e}>3000$ range, while Vicuna and Llama-3.1 showed no improvement. As shown in figure (b), after Knowledge-tuning, Qwen-2.5 and Llama3.1’s performance decreased for apporoximately 33% in the $Rank_{e}>3000$ range compared to Zero-shot. The comprehensive experimental results across all nine encoder-language model combinations are shown in Tab. 2.
Figure 6: Overview of our proposed VisPRE framework. Our framework enriches the vision encoder with entity-specific prior knowledge by first generating (image, caption) pairs from entity text and then finetuning the vision encoder using contrastive loss. The enhanced encoder is subsequently integrated into the MLLM, which is further fine-tuned on Knowledgetuning data.
3.4 Vision Prior Remediation
Figure 7: Improvements of our training framework compared to Zero-shot. With the CLIP encoder, all three models show significant performance gains across different $Rank_{e}$ entities, particularly for those with low vision priors.
In previous sections, we established that MLLM performance correlates positively with vision prior knowledge, and that end-to-end Finetuning yields insufficient. Based on these findings, we propose VisPRE, a training framework that injects entity-related prior knowledge at vision encoder level to enhance MLLM performance. The specific process of our training framework is illustrated in Fig. 6, which comprises two key stages:
Vision Encoder LLM Zero-shot Perception Knowledge Knowledge* Mix* VisPRE(Ours)
OpenAI ViT-L-14 Vicuna-7B 51.22 49.91 54.05 53.48 55.37 56.31
Llama3.1-8B 37.82 39.26 45.67 45.99 44.71 48.24
Qwen2.5-7B 46.05 48.84 54.57 56.59 53.49 54.42
SigLIP ViT-SO-14 Vicuna-7B 52.03 53.66 53.66 57.24 57.07 57.89
Llama3.1-8B 38.91 37.66 41.28 41.84 41.42 41.28
Qwen2.5-7B 36.45 36.31 41.13 41.42 42.84 44.54
DFN ViT-H-14 Vicuna-7B 59.07 58.70 63.33 64.97 62.90 65.56
Llama3.1-8B 38.70 39.84 45.08 46.99 45.69 46.99
Qwen2.5-7B 40.45 38.10 43.33 44.66 46.76 45.40
Table 2: Results on 9 MLLM combinations. Our method outperforms finetuning approaches including Perception-tuning, Knowledge-tuning, Knowledge-tuning* and Mix-tuning*, demonstrating that our method significantly enhances MLLM performance through prior remediation. We mark the best result in bold for each model, and * indicates unfreezing the vision encoder parameters during fine-tuning.
  • Remedy Encoder: We first generate image captions from entity text to form (image, caption) pairs, then finetune the vision encoder alongside the text encoder using contrastive loss.
  • Instruction Tuning: We subsequently integrate the finetuned encoder into the MLLM architecture and finetune the entire model using Knowledge-tuning data.
To systematically evaluate our proposed framework, we establish several baselines: Zero-shot, Perception-tuning, and Knowledge-tuning from Section 3.2. Additionally, we include Knowledge-tuning* and Mix-tuning*, where the asterisk (*) denotes unfreezing the vision encoder parameters during fine-tuning. Mix-tuning represents a combination of Knowledge-tuning and Perception-tuning data. The evaluation results are presented in Tab. 2.
Finding 3: Remediating prior knowledge at vision encoder level is effective. First, Perception-tuning demonstrates only marginal improvements over Zero-shot performance, and in some configurations, performance even degrades. Knowledge-tuning improvements remain limited. This evidence suggests that fine-tuning with frozen encoder parameters fails to effectively enhance MLLM capabilities. Furthermore, Knowledge-tuning* shows modest performance gains over standard Knowledge-tuning. Finally, our VisPRE framework outperforms all baseline approaches, demonstrating substantial improvement and achieving superior results in six out of nine model combinations. These findings indicate that enhancing the encoder’s prior knowledge substantially expands the capability boundaries of MLLMs.
4 In-depth Analysis
In this section, we conduct an in-depth analysis of the reason behind our method’s effectiveness. In Section 4.1, We analyzed two possible causes: gradient diminishment and loss misalignment. In Section 4.2, we provide a case study.
4.1 Unfreezing vs. Contrastive Learning
Notably, although Knowledge-tuning* adjusted the encoder’s parameters during fine-tuning, it did not show a performance advantage over Knowledge-tuning and was outperformed by our method, which first finetune the encoder with contrastive loss before fine-tuning the MLLM. There are two possible explanations:
Vision Encoder LLM Freeze LLM All parameters
SigLIP ViT-SO-14 Vicuna-7B 29.40 52.04
Llama3.1-8B 33.47 36.58
Qwen2.5-7B 48.31 33.89
Table 3: Comparison of Freeze LLM and All parameters Strategy
Gradient Diminishment:
The gradients propagating to the encoder layer may diminish significantly, limiting their effectiveness in adjusting the encoder’s parameters. To validate the impact of gradient propagation, we conducted experiments using two distinct fine-tuning configurations. In the first configuration Freeze LLM, we kept the language model parameters fixed while finetuning only the adapter and vision encoder. In the second configuration All Parameters, we finetune all model components, including the language model, adapter, and vision encoder. Both configurations were trained on a specially constructed Perception-tuning dataset. Details of the dataset are shown in Appendix A.
The experimental results across three model combinations are presented in Tab. 3. Models fine-tuned with All Parameters consistently demonstrated superior performance compared to their Freeze LLM counterparts, showing significant improvements in four model combinations. Notably, even though Freeze LLM ensures strong gradient flow to the vision encoder due to fewer trainable layers, it still underperformed relative to All Parameters. This suggests that diminished gradients are not the primary factor limiting the effectiveness of fine-tuning with frozen parameters.
Vision Encoder LLM Genrative Contrastive
SigLIP ViT-SO-14 Vicuna-7B 52.04 50.94
Llama3.1-8B 36.58 36.99
Qwen2.5-7B 33.89 34.04
Table 4: Comparison of Generative and Contrastive Strategy
Loss Misalignment:
The language model’s predict-next-token loss may not effectively fine-tune the vision encoder, whereas contrastive loss does. To validate the Loss Misalignment hypothesis, we conducted experiments using two datasets: Generative and Contrastive. The Generative dataset is identical to Perception-tuning and serves to evaluate the effectiveness of predict-next-token loss. For the Contrastive dataset, we drew inspiration from CLIP’s contrastive learning methodology to create multiple-choice questions matching Perception-tuning’s size. In this dataset, models must select the correct entity text from several candidates that best matches the given image, allowing us to assess the effectiveness of contrastive loss. During fine-tuning on both datasets, we unfroze the encoder parameters. The results of these experiments are presented in Tab. 4. Our findings reveal that models trained with the Contrastive approach consistently outperform those trained with the Generative approach, indicating that contrastive training is more effective for knowledge acquisition in CLIP. These results provide support for our Loss Misalignment hypothesis.
4.2 Case Study
Figure 8: Examples of Vicuna-7b’s responses with different encoders. When prompted with image description, the LLM answers correctly, demonstrating adequate knowledge of image entities. However, the original (Origin) and fine-tuning with Knowledge-tuning data (SFT) MLLM fails to answer, highlighting the limitations of its vision encoder. With VisPRE(Ours*), the model answer accuratly. For additional cases, refer to Appendix C.
Here we present an illustrative example. As shown in the upper left of Fig. 8, we input an image of the Portuguese Synagogue with the entity-related question: “Where were this synagogue’s books sent in 1979?”. For (1) LLM: The MLLM correctly answers when receiving only the textual description This is Portuguese Synagogue” instead of the actual image, indicating the LLM component possesses knowledge about this entity. For (2) MLLM (Original): With image input, the MLLM fails to answer correctly. We calculated this entity’s $Rank_{e}$ as 516, indicating low prior knowledge in the visual encoder. (3) MLLM (SFT), despite end-to-end fine-tuning, still fails since the visual encoder’s prior knowledge remains unchanged. Our training framework VisPRE first injects prior knowledge into the visual encoder, elevating the entity’s $Rank_{e}$ to 10, then conducting end-to-end fine-tuning. Consequently, (4) Ours* overcomes the visual encoder’s limitations and correctly answers the question.
5 Related Works
Multi-modal Large Language Models.
MLLMs incorporate visual features into language models, enabling them to perform a wide range of visual tasks. The current MLLM implementations can be primarily classified into two categories. (1) Monolithic MLLMs. Tokenizing different modal inputs uniformly and train the model from scratch[34, 5, 9, 41], which is computationally expensive. (2) Modular MLLMs. Utilizing pre-trained vision-language models (e.g., CLIP[33], SigLIP[40], DINOv2[31]) to obtain visual representations of images, and then train MLLMs through cross-modal data, aligning the visual features provided by vision encder to language model’s embedding space. This method is more data-efficient and widely used by open-source MLLMs (e.g., Flamingo[2], BLIP2[20], LLaVA[26], Qwen-VL[4]). Our work focuses on modular multimodal models. While most works treat modular MLLM as a unified system, our research focuses on the impact of vision encoder part on the language model part.
Cross-modality Alignment.
With the increasing adoption of vision encoder integrated MLLMs, there are researches focusing on the relationship between vision encoder and MLLM performance. Tong et al. [37] found that CLIP[33] and corresponding MLLMs have similar performance trends across various visual modalities, indicating that CLIP features are the main cause of MLLM deficiencies in these visual modes, and addressed these deficiencies by introducing DINOv2[31] features. The recent Yang et al. [39] proposed cross-modal alignment and correspondence metrics to measure vision encoder performance, fitting a binary quadratic polynomial that accurately predicts the performance of MLLMs using that vision encoder. Different from previous works, our work provides a new perspective, showing that from a entity-related prior knowledge standpoint, CLIP’s knowledge is positive correlated with the corresponding MLLM’s performance.
6 Conclusion
In this paper, we introduce $Rank_{e}$ to quantify prior knowledge in vision encoder. We find that MLLM’s performance positively correlated with prior knowledge of vision encoder, and end-to-end finetuning MLLM yields insufficient on improving low prior entity performance. To address this issue, we propose VisPRE training framework which enhances MLLM’s performance by increasing the prior knowledge within the vision encoder. Our study demonstrates a novel pathway for enhancing MLLM performance, offering substantial value for applications involving uncommon entities.
A Datasets
Here we describe the detailed construction process of our dataset. Based on Encyclopedia-VQA[28], we constructed Knowledge-tuning and Perception-tuning datasets for each encoder-language model combination to validate Finding 2. Moreover, we developed Generative and Contrastive dataset to examine the Loss Misalignment hypothesis proposed in 4.2.
A.1 Preprocess
Question Filtering. First, we focus on improving the parts where MLLM’s capabilities are limited by the vision encoder. Therefore, we only retained questions that could be answered by the corresponding LLM when prompted with ”This is {entity_name}” instead of the actual image. Next, to ensure that there were no duplicate or similar questions for the same entity across training and test sets, we deduplicated the dataset based on (entity_name, answer) pairs. Finally, we only retained entities with three or more corresponding questions to ensure sufficient questions for dividing into training and validation sets.
Prior Calculation. We calculated $Rank_{e}$ for all entities in the filtered dataset. We examined the distribution of $Rank_{e}$ calculated using different types of encoders (CLIP[33], SigLIP[40], DFN[15]) across the dataset, as shown in Fig. 10. We found significant variations in $Rank_{e}$ distributions among different encoders. CLIP’s $Rank_{e}$ values were mostly concentrated in the range of $Rank_{e}<400$, with entity counts increasing as $Rank_{e}$ decreased; In contrast, SigLIP’s $Rank_{e}$ distribution is more uniform, with at least 10 entities present across most $Rank_{e}$ intervals; DFN’s $Rank_{e}$ distribution was similar to CLIP’s, with most values concentrated in the range of $Rank_{e}<400$.
Entity Sampling. For SigLIP, we divided $Rank_{e}$ into intervals of size 1000 and sampled 10 entities from each interval. For CLIP and DFN, using the same sampling strategy as SigLIP would result in insufficient sampling of entities in dense intervals, making it difficult to distinguish different levels of prior knowledge in these regions. Therefore, we adopted a sampling method that approximates the original distributions of CLIP and DFN. We sampled 10 entities from intervals of $20<Rank_{e}<=$ 2, $42<Rank_{e}<=$ 4, $84<Rank_{e}<=$ 8, …, $1024512<Rank_{e}<=$ 1024, $Rank_{e}>1024$, ensuring that the sampled distribution approximates the original distribution while retaining all entities with low prior knowledge to reflect the relationship between entity prior knowledge and model performance. Finally, we retained the questions corresponding to the sampled entities and divided the dataset into training and test sets, with statistical information shown in Tab. 1.
Figure 9: Knowledge-tuning and Perception-tuning datasets
A.2 Construction
For Knowledge-tuning dataset, we use the original question and answer from the Encyclopedia-VQA dataset. For Perception-tuning dataset, we replace the original question in the Knowledge-tuning dataset with cognitive question like ”What is this image about?” and substitute the answers with the entity text corresponding to the image. Examples of Knowledge-tuning and Perception-tuning datasets are shown in Fig. 9.
For the Generative dataset, we use the same question and answer as in the Perception-tuning dataset but adopt a different prompt template to align with the template used in Contrasting dataset. For Contrasive dataset, instead of directly asking the model to identify the entity in the image, we provide two possible options and ask the model to choose one. We randomly adopt two types of prompts: the first is ”What is this image about?”, and the second is ”Which of the following is NOT in the image?”. For the latter, the model needs to identify the object not present in the image. The prompt templates for the Generative and Contrasive datasets are shown in Fig. 11 and Fig. 12, respectively.
B Evaluation Settings
We employ Llama-3.1-70B[14] to evaluate the accuracy of MLLM’s responses to VQA questions. Specifically, we provide Llama-3.1-70B with the question, entity name (wikipedia_title in prompt), ground truth answer, and MLLM’s response. The model outputs true to indicate a correct answer and false to indicate an incorrect answer. The prompt template is shown in Fig. 13, with the few_shot_examples shown in Fig. 14.
C More Cases
In Fig. 15, we demonstrated Vicuna-7B’s responses under different encoder configurations. Here in Fig. 15, we show examples of responses from Llama-3.1-7B and Qwen-2.5-7B under different encoders.
(a) Vicuna
(b) Qwen-2.5
(c) Llama-3.1
Figure 10: The $Rank_{e}$ distribution of entities calculated using three different encoders. Here we show the entities that (a)Vicuna, (b)Qwen-2.5 and (c)Llama-3.1 could answer after using text prompts instead of entity images. We can see that the $Rank_{e}$ distributions for both CLIP and DFN are concentrated in intervals near the left side, while SigLIP’s $Rank_{e}$ distribution is relatively uniform.
Prompt for Generative dataset You are given an image. Identify what the image is about with one word or short phrase. Example: [Here is an image of Baron Hotel.] Question: What is this image about? Answer: Baron Hotel Now we begin: {entity_image} Question: What is this image about? Answer:
Figure 11: Prompt for Generative dataset. The placeholder entity_image represents the input image
Prompt for Contrasive dataset You are given an image and two options. Your task is to answer based on one of the following two types of questions: Identify what the image is about by selecting one word or short phrase from the options. Identify which of the two options is NOT present in the image. Example: [Here is an image of Baron Hotel.] Question: What is this image about? Options: (A) Markenfield Hall (B) Baron Hotel Answer: Baron Hotel [Here is an image of Baron Hotel.] Question: Which of the following is NOT in the image? Options: (A) Markenfield Hall (B) Baron Hotel Answer: Markenfield Hall Now we begin: {entity_image} Question: {question} Options: (A) {entityA} (B) {entityB} Answer:
Figure 12: Prompt for Contrasive dataset. The placeholder entity_image represents the input image, while question is selected from either ”What is this image about?” or ”Which of the following is NOT in the image?”. entityA and entityB include one correct entity corresponding to the image and one incorrect entity randomly sampled from the dataset.
Prompt for Llama-3.1 evaluation You are an expert evaluator tasked with assessing the correctness of model predictions. Your job is to determine if a given prediction is correct based on the provided information. Follow these strict guidelines: 1. You will be given four pieces of information: - Question: The original question asked - Wikipedia_title: The title of the Wikipedia article that corresponds to the knowledge base for the question - Answer: The correct answer(s) to the question, possibly including multiple candidates separated by ”|||| - Prediction: The model’s prediction to be evaluated 2. Understand that the question is specifically about the entity described in the Wikipedia_title. 3. Compare the prediction to the answer(s), taking into account the context of the question and the Wikipedia_title. 4. Apply these strict criteria: - The prediction must be accurate and specific. - If there are multiple candidate answers separated by ”||||”, the prediction must match at least one of them to be considered true. - For numerical answers, the prediction must be within 10% of at least one correct answer to be considered true. - For categorical or descriptive answers, the prediction must match the key concepts or categories in at least one of the provided answers. - Partial or vague answers that don’t fully capture the specificity of any correct answer should be considered false. - Pay close attention to units, specificity, and context provided in the question, Wikipedia_title, and answer(s). 5. Your response must be exactly one word: - Output ”true” if the prediction meets all the criteria for correctness. - Output ”false” if the prediction fails to meet any of the criteria. 6. Do not provide any explanations or additional comments. {few_shot_examples} Remember, your task is to evaluate the correctness of the prediction based on all the information provided. Be strict in your assessment, but consider all given correct answers. Respond only with ”true” or ”false”. Question: {question} Wikipedia_title: {wikipedia_title} Answer: {answer} Prediction: {prediction} Evaluation:
Figure 13: Complete prompt for evaluating MLLM responses using Llama-3.1-70B. We prompt the model to determine whether a prediction is correct by examining the question, wikipedia_title (entity name), and answer. The model outputs true for correct predictions and false for incorrect ones. The few_shot_examples are shown in Fig. 14
Few-shot examples Examples: Question: Along with the mojave desert, in what desert is this plant found? Wikipedia_title: Acmispon rigidus Answer: Sonoran Desert Prediction: Sonoran Evaluation: true Question: How many people can this stadium host? Wikipedia_title: Mercedes-Benz Stadium Answer: 71,000 |||| 75,000 Prediction: 73,000 Evaluation: true Question: When was this novel first published? Wikipedia_title: To Kill a Mockingbird Answer: 1960 Prediction: 1962 Evaluation: false
Figure 14: few_shot_examples in prompt for Llama-3.1 evaluation. We provide three examples to help the model understand the evaluation requirements.
Figure 15: We present examples of Llama-3.1 and Qwen-2.5’s responses under three encoder setups. When prompted with text to identify objects in the image, the LLM provides correct answers, demonstrating its knowledge of image entities. In contrast, the MLLM (Origin) fails to respond correctly, highlighting the limitations of its vision encoder. Even after fine-tuning with Knowledge-type VQA data (MLLM SFT), the model still cannot provide accurate answers, revealing the constraints of fine-tuning. Finally, with our Remedy Encoder, the model delivers accurate responses, demonstrating that our method effectively expands the MLLM’s visual priors.